# Two-grid convergence theory for symmetric positive semidefinite linear systems

### v1

@@ Abstract.
This paper is devoted to the convergence theory of two-grid methods for symmetric positive semidefinite linear systems, with particular focus on the singular case. In the case where the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver, we derive a succinct identity for characterizing the convergence factor of two-grid methods. More generally, we present some convergence estimates for two-grid methods with approximate coarse solvers, including both linear and general cases. A key feature of our analysis is that it does not require any additional assumptions on the system matrix, especially on its null space.
@@
two-grid methods, symmetric positive semidefinite matrices, singular systems, convergence analysis
@@ 1. Introduction
As a powerful solver with linear or near-linear computational complexity, multigrid has been widely used for the numerical solution of partial differential equations; see, e.g., [11, 4, 26, 27, 31]. The foundation of multigrid is a two-grid scheme, which consists of two complementary error-reduction processes: smoothing (or relaxation) and coarse-grid correction. In the correction step, a coarse-grid system with fewer unknowns needs to be solved. In practice, it is often too costly to solve the coarse-grid system exactly. A typical strategy is to apply the two-grid scheme recursively in the correction steps, which will yield an algorithm with multilevel structure (i.e., multigrid method). For symmetric positive definite (SPD) linear systems, the convergence properties of two-grid and multigrid methods have been well studied; see, e.g., [29, 35, 30, 7, 8, 17, 31, 32, 33, 34].
@@
Many systems arising in mathematical models may only be semidefinite, such as discretized partial differential equations with Neumann boundary conditions (see, e.g., [1]) and Laplacian matrices of graphs (see, e.g., [24, 16]). In such cases, the system matrix, denoted by $A$, is singular and symmetric positive semidefinite (SPSD). Some convergence properties of classical iterative methods for SPSD systems can be found, e.g., in [13, 5, 14, 6, 9, 28, 10]. The convergence of Schwarz methods for SPSD systems was studied in [21]. Our focus here is on the convergence of two-grid methods for SPSD systems. In practice, two-grid convergence theory can be used to guide or motivate the design of multigrid methods; see, e.g., [3, 19, 18]. Moreover, two-grid convergence estimates can be applied to the analysis of multigrid methods; see, e.g., [33, 34].
@@
In the SPD case, an identity for the convergence factor of two-grid methods was established in [8]; see [30, 36] for an abstract version. Ideally, one can derive a two-grid convergence identity for the semidefinite case without imposing any restrictions on the null space of $A$, which is denoted by $\mathcal{N}(A)$. Under several essential assumptions on local null spaces, a sharp convergence estimate for the successive subspace correction methods [29, 30] applied to SPSD systems was presented in [15]; see [28] for a simplified analysis. More specifically, [15] offers an identity for the seminorm of error propagation operator, despite it involves a ‘sup-inf-inf’ expression relating to the global null space and its orthogonal complement. In [2], a classical two-grid convergence estimate for SPD systems [20, 23, 25] was extended to the semidefinite case. A two-grid convergence theory for singular systems was developed in [22], in which a restrictive condition on the coarse space and $\mathcal{N}(A)$ is required. Even in the special SPSD case, the two-grid estimates in [22] involve an extended prolongation operator, whose construction is based on $\mathcal{N}(A)$. A noteworthy contribution is the two-grid convergence identity appearing in [31, Theorem 5.3], which was established under the condition that $\mathcal{N}(A)$ is contained in the coarse space.
The aforementioned two-grid convergence theory and the associated assumptions rely largely on $\mathcal{N}(A)$, while the properties of $\mathcal{N}(A)$ may not be easily available. Consequently, it might be tough to apply those results or check the required conditions. Motivated by this observation, we establish a new convergence theory for two-grid methods applied to SPSD systems, which lifts the restrictions on $\mathcal{N}(A)$. Our main contribution is twofold.
@@
  • Firstly, we derive a succinct identity for the convergence factor of two-grid methods, provided that the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver (it is called the exact coarse solver).
  • Secondly, we present some convergence estimates for two-grid methods with inexact coarse solvers, including both linear and general cases.
@@
  • It does not require any additional assumptions on $A$, especially on $\mathcal{N}(A)$.
  • It yields an identity for the convergence factor of exact two-grid methods instead of an upper bound.
  • It offers some convergence bounds for two-grid methods with inexact coarse solvers.
The rest of this paper is organized as follows. In Section 2, we give a sufficient and necessary condition for two-grid convergence factor to be less than $1$. In Section 3, we present a concise identity for the convergence factor of exact two-grid methods. In Section 4, we study the convergence of two-grid methods with approximate coarse solvers. In Section 5, we make some concluding remarks.
@@ 2. Preliminaries
  • For two symmetric matrices $S_{1}$ and $S_{2}$, the symbol $S_{1}\succ S_{2}$ (resp., $S_{1}\succeq S_{2}$) means that $S_{1}-S_{2}$ is positive definite (resp., positive semidefinite).
@@
  • $\mathcal{R}(\cdot)$ denotes the range (or column space) of a matrix.
  • $\mathcal{N}(\cdot)$ denotes the null space of a matrix.
  • $\lambda_{\min}(\cdot)$ and $\lambda_{\max}(\cdot)$ denote the smallest and largest eigenvalues of a matrix, respectively.
  • $\lambda_{i}(\cdot)$ denotes the $i$th smallest eigenvalue of a matrix.
@@
The following lemma gives two simple but useful properties of the $A$-seminorm.
@@
(2.3) $\|X\|_{A}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}}.$
@@
It can be readily seen from (2.1) that (2.2) holds for $\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)$. If $\mathbf{v}\in\mathcal{N}(A)$, then
@@
$=\sup_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}},$
@@
Consider the solution of linear systems of the form
@@
where $A\in\mathbb{R}^{n\times n}\backslash\{0\}$, $A\succeq 0$, $\mathbf{u}\in\mathbb{R}^{n}$, and $\mathbf{f}\in\mathcal{R}(A)$. When $A$ is singular (namely, $\mathcal{N}(A)\neq\{0\}$), the system (2.4) has infinitely many solutions for any $\mathbf{f}\in\mathcal{R}(A)$. We aim to solve (2.4) using a two-grid method, the output of which is an approximation to a solution to (2.4). Two very basic assumptions involved in the two-grid analysis are made below.
@@
We remark that $M$ is not necessarily nonsingular. Moreover, measured in $\|\cdot\|_{A}$, the smoothing iteration matrix in Algorithm 1 below (i.e., $I-MA$) is not required to be a strict contraction. Due to
@@
$\|I-MA\|_{A}^{2}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\big{\|}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\big{\|}_{2}^{2}}{\|\mathbf{v}\|_{2}^{2}}=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}},$
@@
Thus, $\|I-MA\|_{A}\leq 1$ if and only if $\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}\geq 0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$, or, equivalently, $\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}\geq 0$ for all $\mathbf{v}\in\mathbb{R}^{n}\backslash\{0\}$ (i.e., $A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\succeq 0$), because $\mathbb{R}^{n}$ is the direct sum of $\mathcal{R}(A)$ and $\mathcal{N}(A)$, denoted by $\mathbb{R}^{n}=\mathcal{R}(A)\oplus\mathcal{N}(A)$.
@@
1:Smoothing: $\mathbf{u}^{(1)}\leftarrow\mathbf{u}^{(0)}+M\big{(}\mathbf{f}-A\mathbf{u}^{(0)}\big{)}$triangleright $\mathbf{u}^{(0)}\in\mathbb{R}^{n}$ is an initial guess
@@
The output of Algorithm 1 converges to a particular solution to (2.4) as $\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}$ tends to zero. Indeed, $\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}=0$ implies that $\mathbf{u}-\mathbf{u}_{\rm TG}\in\mathcal{N}(A)$ and hence $A\mathbf{u}_{\rm TG}=A\mathbf{u}=\mathbf{f}$.
@@
$\Pi^{T}=\Pi=\Pi^{2},\quad\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)},\quad\text{and}\quad\mathcal{R}(\Pi)=\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}.$
Thus, $\Pi$ is an $L^{2}$-orthogonal projector along (or parallel to) $\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}$ onto $\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}$. Since
@@
$\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}\leq\|E_{\rm TG}\|_{A}\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A},$
@@
$\|E_{\rm TG}\|_{A}^{2}$ $=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
$=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F_{\rm TG}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}},$
@@
Hence, it holds that $\|E_{\rm TG}\|_{A}<1$ if and only if $\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$.
The following lemma provides a sufficient and necessary condition (namely, (2.10)) for $\|E_{\rm TG}\|_{A}<1$. It is worth pointing out that (2.10) is not really a restriction on $\mathcal{N}(A)$. In view of (2.10), a sufficient condition for $\|E_{\rm TG}\|_{A}<1$ is $\overline{M}\succ 0$, which in fact can be satisfied by the most commonly used weighted Jacobi and Gauss–Seidel smoothers; see Remark 2.6. Throughout this paper, we always assume that (2.10) holds.
@@
Under the assumptions of Algorithm 1, $\|E_{\rm TG}\|_{A}<1$ if and only if
(2.10) $\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{N}\big{(}P^{T}(I-AM)A^{\frac{1}{2}}\big{)}=\mathcal{N}(A).$
@@
$\mathbf{v}^{T}F_{\rm TG}\mathbf{v}=\underbrace{\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}}_{\geq 0}+\underbrace{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}_{\geq 0}\geq 0\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}$
@@
which contradicts with the condition $\|E_{\rm TG}\|_{A}<1$, or, equivalently, $\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$.
(2.10)$\Rightarrow\|E_{\rm TG}\|_{A}<1$”: Conversely, if (2.10) holds, then $\|E_{\rm TG}\|_{A}<1$. Otherwise, there must exist a vector $\mathbf{y}\in\mathcal{R}(A)\backslash\{0\}$ such that $\mathbf{y}^{T}F_{\rm TG}\mathbf{y}=0$. Then
@@
Under the assumptions of Algorithm 1, if $\overline{M}\succeq 0$ and
@@
From Corollary 2.5, we deduce that $\|E_{\rm TG}\|_{A}<1$ if $\overline{M}\succ 0$, which can be satisfied by the weighted Jacobi and Gauss–Seidel smoothers, provided that the diagonal entries of $A$ are all nonzero. Let
@@
where $D$ and $L$ denote the diagonal and strictly lower triangular parts of $A$, respectively. Without loss of generality, we assume that $D\succ 0$. If some diagonal entry of $D$ is zero, the corresponding row and column in $A$ are both zero, in which case one may solve a reduced version of (2.4) obtained by deleting the zero row and column. If $M=\omega D^{-1}$ with $0<\omega<\frac{2}{\lambda_{\max}(D^{-1}A)}$, then
@@
Another typical choice is $M=(D+L)^{-1}$, in which case one has

### v2

# Convergence analysis of two-grid methods for symmetric positive semidefinite systems

@@ Abstract.
Two-grid theory plays a fundamental role in the design and analysis of multigrid methods. This paper is devoted to a new convergence analysis of two-grid methods for singular and symmetric positive semidefinite systems. Specifically, we derive a concise identity for characterizing the convergence factor of two-grid methods, with the Moore–Penrose inverse of coarse-grid matrix being used as a coarse solver. Furthermore, we present a convergence estimate for two-grid methods with approximate coarse solvers. Our new theory does not require any additional assumptions on the coefficient matrix, especially on its null space.
@@
multigrid methods, two-grid methods, symmetric positive semidefinite matrices, null space, Moore–Penrose inverse
@@ 1. Introduction
Multigrid is a powerful solver with linear or near-linear computational complexity that has been widely used for the numerical solution of partial differential equations; see, e.g., [12, 5, 30, 31, 36]. The foundation of multigrid is a two-grid scheme, which consists of two complementary error-reduction processes: smoothing (or relaxation) and coarse-grid correction. In general, the smoothing process is chosen as a simple iterative method (e.g., the weighted Jacobi and Gauss–Seidel methods), while these classical methods are commonly only effective at reducing high-frequency (or oscillatory) error. The remaining low-frequency (or smooth) error will be treated by the correction process, in which a coarse-grid system with fewer unknowns needs to be solved. In practice, it is often too costly to solve the coarse-grid system exactly. A typical strategy is to apply the two-grid scheme recursively in the correction steps. The resulting methods (namely, multigrid methods) possess a multilevel structure.
@@
For symmetric positive definite (SPD) linear systems, the convergence of multigrid methods has been well studied; see, e.g., [33, 41, 34, 8, 9, 17, 36, 38, 39, 40, 37]. Many important models or methods in mathematics may lead to symmetric positive semidefinite (SPSD) linear systems, such as partial differential equations with pure Neumann boundary conditions (see, e.g., [1]), generalized finite element methods (see, e.g., [26, 27, 28, 35]), and Laplacian matrices of graphs (see, e.g., [21, 25, 16]). Convergence of additive and multiplicative Schwarz methods for SPSD systems was studied in [22]. Convergence theory of general stationary iterative methods for such systems can be found, e.g., in [13, 6, 14, 7, 10, 32, 11].
Our focus here is on two-grid methods for SPSD systems. Two-grid theory plays a fundamental role in the design or analysis of multigrid methods; see, e.g., [3, 19, 18, 39, 40, 37]. In the SPD case, an elegant identity for characterizing the convergence factor of two-grid methods was established in [9]; see [34, 42] for an abstract version. The identity is a powerful tool for analyzing the properties of two-grid methods; see, e.g., [9, 4, 38]. Similarly, we expect to establish a convergence identity for two-grid methods applied to SPSD systems, without requiring any additional conditions on the null space of the coefficient matrix $A$, denoted by $\mathcal{N}(A)$.
@@
Under several essential conditions on local null spaces, a sharp convergence estimate for successive subspace correction methods [33, 34] applied to SPSD systems was presented in [15]; see [32] for a simplified analysis. More specifically, [15] offers an identity for the seminorm of error propagation operator, which involves a ‘sup-inf-inf’ expression relating to the global null space and its orthogonal complement. In [2], a classical two-grid convergence estimate for SPD systems (see [20, 24, 29]) was extended to the semidefinite case. A two-grid convergence theory for singular systems was developed in [23], in which a condition on the coarse space and $\mathcal{N}(A)$ is required. Moreover, an extended prolongation operator, constructed from $\mathcal{N}(A)$, is needed. A noteworthy contribution is the two-grid identity in [36, Theorem 5.3], which was proved under the condition that $\mathcal{N}(A)$ is contained in the coarse space.
In some cases, however, the full information of $\mathcal{N}(A)$ may not be readily available. For example, as indicated in [26, page 52], the eigenfunctions associated with the zero eigenvalues are generally unknown in generalized finite element methods. In addition, most of the existing results focus on the case where the coarse-grid system is solved exactly, in the sense that a generalized inverse of the coarse-grid matrix is used as a coarse solver. For singular systems, more general coarse solvers are seldom considered in the literature. The difficulty may arise from the fact that the error propagation operator of coarse-grid correction fails to be a projector when the coarse-grid system is solved approximately.
In this paper, we establish a new two-grid convergence theory for SPSD systems. Our main contribution is twofold:
@@
  • We derive a succinct identity for the convergence factor of two-grid methods, with the Moore–Penrose inverse of coarse-grid matrix being used as a coarse solver (for convenience, it is called an exact coarse solver);
  • We present a convergence estimate for two-grid methods with inexact coarse solvers (various approximation strategies can be utilized to solve the coarse-grid system).
@@
  • It does not require any additional conditions on the coefficient matrix, especially on its null space.
  • It gives an identity for the convergence factor of exact two-grid methods instead of an upper bound.
  • It provides a convergence estimate for two-grid methods with inexact coarse solvers, that is, our analysis is not confined to the exact case.
The rest of this paper is organized as follows. In Section 2, we give a sufficient and necessary condition for two-grid convergence factor to be less than $1$. In Section 3, we present a concise identity for the convergence factor of exact two-grid methods. In Section 4, we study the convergence of two-grid methods with approximate coarse solvers, and provide a numerical example to illustrate our theoretical estimate. In Section 5, we make some concluding remarks.
@@ 2. Preliminaries
  • For a symmetric matrix $S$, $S\succ 0$ and $S\succeq 0$ mean that $S$ is positive definite and positive semidefinite, respectively.
@@
  • $\mathcal{R}(\cdot)$ denotes the column space (or range) of a matrix.
  • $\mathcal{N}(\cdot)$ denotes the null space (or kernel) of a matrix.
  • $\lambda_{\min}(\cdot)$ and $\lambda_{\max}(\cdot)$ denote the smallest and largest eigenvalues of a matrix with real eigenvalues, respectively.
  • $\lambda_{i}(\cdot)$ denotes the $i$th smallest eigenvalue of a matrix with real eigenvalues.
@@
The following lemma gives two elementary but useful properties of the $A$-seminorm.
@@
[…]
@@
It can be readily seen from (2.1) that the inequality (2.2) holds for $\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)$. If $\mathbf{v}\in\mathcal{N}(A)$, then
@@
[…]
@@
Consider solving the linear system
@@
where $A\in\mathbb{R}^{n\times n}\backslash\{0\}$ is SPSD, $\mathbf{u}\in\mathbb{R}^{n}$, and $\mathbf{f}\in\mathcal{R}(A)$ (that is, $\mathbf{f}$ is orthogonal to $\mathcal{N}(A)$). When $A$ is singular (i.e., $\mathcal{N}(A)\neq\{0\}$), the system (2.4) has infinitely many solutions for any $\mathbf{f}\in\mathcal{R}(A)$. We will analyze the convergence of two-grid methods for solving (2.4). Two basic assumptions involved in the analysis of two-grid methods are made below.
  • (A1):
@@
In view of (A1) and (A2), $M$ is not necessarily nonsingular, and $P$ is not necessarily of full column rank. In addition, measured in $\|\cdot\|_{A}$, the smoothing iteration in Algorithm 1 below is not required to be a strict contraction. Due to
@@
[…]
@@
[…]
@@
1:Smoothing: $\mathbf{u}^{(1)}\leftarrow\mathbf{u}^{(0)}+M\big(\mathbf{f}-A\mathbf{u}^{(0)}\big)$ $\triangleright$ $\mathbf{u}^{(0)}\in\mathbb{R}^{n}$ is an initial guess
@@
In Algorithm 1, the coarse-grid system to be solved reads
(2.6) $A_{\rm c}\mathbf{x}_{\rm c}=\mathbf{r}_{\rm c}.$
It is easy to see that $\mathbf{e}_{\rm c}=A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$ is a solution to (2.6), because $A_{\rm c}A_{\rm c}^{\dagger}$ is a projector onto $\mathcal{R}(A_{\rm c})$ and
$\mathbf{r}_{\rm c}=P^{T}\big(\mathbf{f}-A\mathbf{u}^{(1)}\big)=P^{T}A\big(\mathbf{u}-\mathbf{u}^{(1)}\big)\in\mathcal{R}\big(P^{T}A^{\frac{1}{2}}\big)=\mathcal{R}(A_{\rm c}).$
@@
$\Pi^{2}=\Pi=\Pi^{T},\quad\mathcal{N}(\Pi)=\mathcal{N}\big(P^{T}A^{\frac{1}{2}}\big),\quad\text{and}\quad\mathcal{R}(\Pi)=\mathcal{R}\big(A^{\frac{1}{2}}P\big),$
i.e., $\Pi$ is an $L^{2}$-orthogonal projector along (or parallel to) $\mathcal{N}\big(P^{T}A^{\frac{1}{2}}\big)$ onto $\mathcal{R}\big(A^{\frac{1}{2}}P\big)$. Since
@@
$\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}\leq\|E_{\rm TG}\|_{A}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A},$
@@
[…]
@@
[…]
Remark 2.4.
If $\|E_{\rm TG}\|_{A}<1$ and Algorithm 1 is carried out $\nu$ times iteratively (the corresponding output is denoted by $\mathbf{u}_{\rm TG}^{(\nu)}$), then $\|\mathbf{u}-\mathbf{u}_{\rm TG}^{(\nu)}\|_{A}\to 0$ as $\nu\to\infty$, in which case $\big\{\mathbf{u}_{\rm TG}^{(\nu)}\big\}_{\nu=1}^{\infty}$ converge to a particular solution to (2.4). Indeed, $\|\mathbf{u}-\mathbf{u}_{\rm TG}^{(\nu)}\|_{A}=0$ implies that $\mathbf{u}-\mathbf{u}_{\rm TG}^{(\nu)}\in\mathcal{N}(A)$ and hence $A\mathbf{u}_{\rm TG}^{(\nu)}=A\mathbf{u}=\mathbf{f}$.
The following lemma provides a sufficient and necessary condition (namely, (2.11)) for $\|E_{\rm TG}\|_{A}<1$. It is worth pointing out that (2.11) is not really a restriction on $\mathcal{N}(A)$. In fact, we can deduce from (2.11) that a sufficient condition for $\|E_{\rm TG}\|_{A}<1$ is $\overline{M}\succ 0$. This condition can be satisfied by the conventional weighted Jacobi and Gauss–Seidel smoothers; see Remark 2.7 below. Throughout this paper, we always assume that (A1), (A2), and (2.11) hold.
@@
The convergence factor $\|E_{\rm TG}\|_{A}$ is strictly less than 11 if and only if
(2.11) $\mathcal{N}\big(A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big)\cap\mathcal{N}\big(P^{T}(I-AM)A^{\frac{1}{2}}\big)=\mathcal{N}(A),$
where $\overline{M}$ is defined by (2.5).
@@
$\mathbf{v}^{T}F_{\rm TG}\mathbf{v}=\underbrace{\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}}_{\geq\,0}+\underbrace{\mathbf{v}^{T}\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\Pi\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)\mathbf{v}}_{\geq\,0}\geq 0\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}$
@@
[…]
@@
Let $\overline{M}$ be defined by (2.5). If $\overline{M}\succeq 0$ and
@@
Let
@@
where $D$ and $L$ denote the diagonal and strictly lower triangular parts of $A$, respectively. Without loss of generality, we assume that $D\succ 0$. If some diagonal entry of $D$ is zero, the corresponding row and column in $A$ are both zero, in which case one may solve a reduced version of (2.4) obtained by deleting the zero row and column. From Corollary 2.6, we deduce that $\|E_{\rm TG}\|_{A}<1$ if $\overline{M}\succ 0$, which can be satisfied by the most commonly used weighted Jacobi and Gauss–Seidel smoothers.
  • If $M=\omega D^{-1}$ with $0<\omega<\frac{2}{\lambda_{\max}(D^{-1}A)}$, then
@@
  • If $M=(D+L)^{-1}$, then